# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v2

@@ VII-J Proof of 4
Assume on the contrary that there exists some $N>0$, such that for every population size $\mathsf{N}\in\mathcal{N}_{\geq N}$, where $\mathcal{N}_{\geq N}=\{\mathsf{N}|\mathsf{N}\in\mathcal{N},\mathsf{N}\geq N\}$, the evolution of the proportion of $\mathtt{A}$-players admits some perpetual fluctuation set ${\mathcal{Y}}^{\mathsf{N}}\subset(0,1)\cap\frac{1}{\mathsf{N}}\mathbb{Z}$ of length $y^{\mathsf{N}}$, where $y^{\mathsf{N}}\geq L$ for some positive value $L$,
@@
and that for every $K\in(0,L]$ there exists some small enough $\delta>0$ such that in the long run the proportion of $\mathtt{A}$-players visits every connected subset of length $K$ of ${\mathcal{Y}}^{\mathsf{N}}$ (for some $K>0$) with probability greater than $\delta$:
@@
where $\mu_{x}^{\frac{1}{\mathsf{N}}}$ is the stationary measure of the Markov chain associated with the evolution of the population proportion of $\mathtt{A}-$players $x^{\mathsf{N}}$ whose existence is straightforward to show. It can be shown that the family of these Markov chains is a GSAP for the abstract dynamics (9), and the Birkhoff center $\mathrm{BC}^{x}$ of the dynamical system induced by the abstract dynamics is the union of all equilibrium points. Accordingly in view of Theorem 1 and similar to how Theorem 3 was proven, it follows that for every vanishing sequence $\langle\frac{1}{n}\rangle=\frac{1}{n_{0}},\frac{1}{n_{1}},\ldots$ where $n_{i}\in\mathcal{N}$ for $i\in\mathbb{Z}_{\geq 0}$ and for every open set $\mathcal{O}$ containing $\mathrm{BC}^{x}$, we have $\lim_{i\to\infty}\mu_{x}^{1/n_{i}}({\mathcal{O}})=1.$ Since the set $\mathrm{BC}^{x}$ consists of isolated points, it can be shown that, starting from a specified initial condition, there exists one member of $\mathrm{BC}^{x}$, say point ${q}$, such that $\lim_{i\to\infty}\mu_{x}^{\frac{1}{n_{i}}}({\mathcal{B}}({q},\epsilon/2))=1$ for small enough $\epsilon$. This means that for every $\delta>0$, there exists $i^{*}>0$ such that for all $i>i^{*}$, $\mu^{\frac{1}{n_{i}}}({\mathcal{B}}({q},\epsilon/2))>1-\delta$, and, in turn, $\mu^{\frac{1}{n_{i}}}([\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-{\mathcal{B}}({q},\epsilon/2))<\delta$. On the other hand, in view of (19), $[\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-{\mathcal{B}}({q},\epsilon/2)$ contains a connected interval of length at least $(L-\epsilon)/2$, which contradicts (20) for $K=(L-\epsilon)/2$. ∎
VII-K 2
@@
As the update mechanism is asynchronous, the population states at two consecutive time indices $\bm{x}^{\mathsf{N}}(k)$ and $\bm{x}^{\mathsf{N}}(k+1)$ can differ by at most ±$\pm\frac{1}{\mathsf{N}}$. Which element of the population state $\bm{x}^{\mathsf{N}}(k+1)$ may differ from that of $\bm{x}^{\mathsf{N}}(k)$ is determined by the random variable ${P}_{k}$ which is the active subpopulation at time index $k$, and can take values in $[\mathsf{p}+\mathsf{p}^{\prime}]$. The probability of ${P}_{k}=p$ equals $\rho_{p}$. The amount of change in the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ depends on the current strategy of the active agent and the preferred strategy of the subpopulation ${P}_{k}$. The former is captured by the random variable ${S}_{k}$ taking values $1$ and $2$, and the latter is determined by the function $s^{*}({P}_{k},\bm{x}^{\mathsf{N}}(k))$ which is obtained from (1) and (2). The probability of ${S}_{k}=1$ (resp. ${S}_{k}=2$) given the active subpopulation is $p$ is $x^{\mathsf{N}}_{p}/\rho_{p}$ (resp. $1-x^{\mathsf{N}}_{p}/\rho_{p}$). If the current strategy and the preferred strategy are equal, the population states at time indices $k$ and $k+1$ are the same, and if the current strategy is $\mathtt{B}$ and the preferred one is $\mathtt{A}$, the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ increases by $\frac{1}{\mathsf{N}}$, and the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ decreases by $\frac{1}{\mathsf{N}}$ otherwise. This update rule is equivalent to (3).

### v3

@@ VII-K Proof of 4
Assume on the contrary that there exists some $N>0$, such that for every population size $\mathsf{N}\in\mathcal{N}_{\geq N}$, where $\mathcal{N}_{\geq N}=\{\mathsf{N}|\mathsf{N}\in\mathcal{N},\mathsf{N}\geq N\}$, the evolution of the proportion of $\mathtt{A}$-players admits some perpetual fluctuation set ${\mathcal{Y}}^{\mathsf{N}}\subset(0,1)\cap\frac{1}{\mathsf{N}}\mathbb{Z}$ of amplitude $y^{\mathsf{N}}$, where $y^{\mathsf{N}}\geq L$ for some positive value $L$,
@@
and that for every $K\in(0,L]$ there exists some small enough $\delta>0$ such that in the long run the proportion of $\mathtt{A}$-players visits every connected subset of amplitude $K$ of ${\mathcal{Y}}^{\mathsf{N}}$ (for some $K>0$) with probability greater than $\delta$:
@@
where $\mu_{x}^{\frac{1}{\mathsf{N}}}$ is defined as $\mu^{\frac{1}{\mathsf{N}}}\circ\bm{f}^{-1}$, where $\bm{f}:\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}\to\mathbb{R},$ and $\bm{f}(\bm{x})=\sum_{p=1}^{\mathsf{p}}x_{p}.$ Let $\mathtt{n}$ denote the total number of abstract equilibrium points, and let $q_{p}$, for $p\in[\mathtt{n}]$, denote the $p^{\text{th}}$ equilibrium point of the abstract dynamics. Given Theorem 3, it follows that for every vanishing sequence $\langle\frac{1}{n}\rangle=\frac{1}{n_{0}},\frac{1}{n_{1}},\ldots$ where $n_{i}\in\mathcal{N}$ for $i\in\mathbb{Z}_{\geq 0}$ $\lim_{i\to\infty}\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))=1,$ for a small enough positive $\epsilon<L/\mathtt{n}$. Equivalently, for every $\delta>0$, there exists $i^{*}>0$ such that for all $i>i^{*}$, $\smash{\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))>1-\delta}$ and, in turn, $\smash{\mu_{x}^{\frac{1}{n_{i}}}([\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))<\delta}$. On the other hand, in view of (20), $[\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2)$ contains a connected interval of amplitude at least $(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$, This contradicts (21) for $K=(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$. ∎
VII-L A3
@@
As the update mechanism is asynchronous, the population states at two consecutive time indices $\bm{x}^{\mathsf{N}}(k)$ and $\bm{x}^{\mathsf{N}}(k+1)$ can differ by at most ±$\pm\frac{1}{\mathsf{N}}$. Which element of the population state $\bm{x}^{\mathsf{N}}(k+1)$ may differ from that of $\bm{x}^{\mathsf{N}}(k)$ is determined by the random variable ${P}_{k}$ which is the active subpopulation at time index $k$, and can take values in $[\mathsf{p}+\mathsf{p}^{\prime}]$. The probability of ${P}_{k}=p$ equals $\rho_{p}$. The amount of change in the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ depends on the current strategy of the active agent and her preferred strategy. The former is captured by the random variable ${S}_{k}$ taking values $1$ and $2$, and the latter is determined by the function $s^{*}({P}_{k},\bm{x}^{\mathsf{N}}(k),{S_{k}})$ which is obtained from (1) and (2). The probability of ${S}_{k}=1$ (resp. ${S}_{k}=2$) given the active subpopulation is $p$ is $x^{\mathsf{N}}_{p}/\rho_{p}$ (resp. $1-x^{\mathsf{N}}_{p}/\rho_{p}$). If the current strategy and the preferred strategy are equal, the population states at time indices $k$ and $k+1$ are the same, and if the current strategy is $\mathtt{B}$ and the preferred one is $\mathtt{A}$, the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ increases by $\frac{1}{\mathsf{N}}$, and the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ decreases by $\frac{1}{\mathsf{N}}$ otherwise. This update rule is equivalent to (3). ∎
Corollary A1.
Assume that an active agent $i$ chooses her strategy uniformly at random when $x_{-i}^{\mathsf{N}}(k)=\tau(i)$. Then, Lemma 1 remains true.
Proof.
It is easy to show that Conditions 1, 2, and 4 in Definition 4 are satisfied. Now, we investigate Condition 3. Function $s^{*}(\cdot,\cdot,\cdot)$ changes to the random variable $S^{*}$ with the following distribution
$\mathbb{P}[$ $S^{*}(p,\bm{x}^{\mathsf{N}},s)=1\mid\bm{x}^{\mathsf{N}}]=$
$\begin{cases}\frac{1}{2}&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p},\text{ and }s=2,\\ &\text{or }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\\ &\text{if }p>\mathsf{p},{x}^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},{x}^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 1&\text{if }p\leq\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau_{p},\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 0&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\end{cases}$
$\mathbb{P}[$ $S^{*}(p,\bm{x}^{\mathsf{N}},s)=2\mid\bm{x}^{\mathsf{N}}]$
$\begin{cases}\frac{1}{2}&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 0&\text{if }p\leq\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau_{p},\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ 1&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1.\\ \end{cases}$
Following the steps taken in the proof of Lemma 1, the expected change in the population state $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$ can be obtained, where for all $\bm{x}\in\bm{\mathcal{X}}_{ss}$ satisfying $\nexists\tau_{p}(\bm{1}^{\top}\bm{x}\in[\tau_{p},\tau_{p}+\frac{1}{\mathsf{N}}])$ is independent of the tie-breaking rule. Out of those states satisfying $\exists\tau_{p}(\bm{1}^{\top}\bm{x}\in[\tau_{p},\tau_{p}+\frac{1}{\mathsf{N}}])$, we investigate the state $\bm{x}$ with $\bm{1}^{\top}\bm{x}=\tau_{p}$ for some $p$; the other values of $\bm{x}$ can be handled similarly. The $p^{\text{th}}$ element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, which corresponds to subpopulation $p$, at $\bm{x}$ with $\bm{1}^{\top}\bm{x}=\tau_{p}$ equals $\mathsf{N}\left(\frac{1}{\mathsf{N}}(\rho_{p}-x_{p})\frac{1}{2}-\frac{1}{\mathsf{N}}x_{p}\times\big{(}0\big{)}\right)=\frac{\rho_{p}}{2}-\frac{x_{p}}{2},$ if $p$ corresponds to an anticoordinating subpopulation and $\mathsf{N}\left(\frac{1}{\mathsf{N}}(\rho_{p}-x_{p})\frac{1}{2}-\frac{1}{\mathsf{N}}x_{p}\big{(}1\big{)}\right)=\frac{\rho_{p}}{2}-\frac{3}{2}x_{p},$ otherwise. In both cases the expected change in subpopulation $p$ belongs to $\mathcal{V}_{p}(\bm{x})=[-x_{p},\rho_{p}-x_{p}]$ (equation (7)). For elements $l\neq p$ of $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, the expected change is equal to the unique member of the $\mathcal{V}_{l}(\bm{x})$. This implies that equation (5) is satisfied with $\bm{y}=\bm{x}$. Hence, Condition 3 holds, and this completes the proof. ∎